OpenStudy (ny,ny):

A ball's height above the ground is modeled by the equation h(t) = -16t^2 + 80t + 40 When does the ball hit the ground?

OpenStudy (baru):

set h(t)=0 and then solve for t

OpenStudy (ny,ny):

I tried, but I got 8(-2t^2+10t+5) = 0 what do I now? @baru

OpenStudy (baru):

\(-2t^2 + 10t + 5=0\) this is a quadratic equation. you have to use the quadratic formula

OpenStudy (ny,ny):

\[t = { -10 \pm \sqrt{140}}\]

OpenStudy (baru):

yes, simplify that, use a calculator

OpenStudy (baru):

also, have you done the quadratic formula correctly? \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (ny,ny):

over -4

OpenStudy (baru):

? i dont get what u mean

OpenStudy (baru):

oh, ok, yes your answer over -4

OpenStudy (ny,ny):

I have to go to class sorry. ty though.

OpenStudy (baru):

sure :)